# Geometric Distortion Guided Transformer for Omnidirectional Image Super-Resolution

### v1

@@ Abstract
As virtual and augmented reality applications gain popularity, omnidirectional image (ODI) super-resolution has become increasingly important. Unlike 2D plain images that are formed on a plane, ODIs are projected onto spherical surfaces. Applying established image super-resolution methods to ODIs, therefore, requires performing equirectangular projection (ERP) to map the ODIs onto a plane. ODI super-resolution needs to take into account geometric distortion resulting from ERP. However, without considering such geometric distortion of ERP images, previous deep-learning-based methods only utilize a limited range of pixels and may easily miss self-similar textures for reconstruction. In this paper, we introduce a novel Geometric Distortion Guided Transformer for Omnidirectional image Super-Resolution (GDGT-OSR). Specifically, a distortion modulated rectangle-window self-attention mechanism, integrated with deformable self-attention, is proposed to better perceive the distortion and thus involve more self-similar textures. Distortion modulation is achieved through a newly devised distortion guidance generator that produces guidance by exploiting the variability of distortion across latitudes. Furthermore, we propose a dynamic feature aggregation scheme to adaptively fuse the features from different self-attention modules. We present extensive experimental results on public datasets and show that the new GDGT-OSR outperforms methods in existing literature.
@@ I Introduction
One way to tackle this issue is through image super-resolution (SR), which reconstructs a high-resolution (HR) image from a low-resolution (LR) input [2, 3]. For convenience in storage and transmission, raw ODIs are generally projected into 2D planar representations. Equirectangular projection (ERP) is the most common method for representing ODIs [4].
@@
Transformer has emerged as a powerful and versatile computational paradigm [6]. Local square-window self-attention is proposed to reduce the computational complexity of the global self-attention mechanism in vision transformer, leading to a limitation in the receptive field [7, 8]. However, for omnidirectional image super-resolution (ODISR), an inconsistency exists between the distortion direction and the shape of the squared windows, as the ERP images are unevenly stretched. For example, a circle in the ODI is distorted into an oval in the ERP image. Small squared windows struggle to capture the whole oval, while large squared windows may encompass irrelevant patterns. Therefore, square-window self-attention is a suboptimal option for reconstructing ERP images. Rectangular windows [8] can calibrate and expand the receptive field by involving more self-similar textures along the direction of stretching distortion, which is more appropriate for modeling ERP images than squared windows. However, existing ODISR methods fail to reconstruct ERP images according to their geometric distortion, leading to the limitation of involved pixels and self-similar textures. To solve this issue, we propose a Geometric Distortion Guided Transformer for Omnidirectional image Super-Resolution (GDGT-OSR), which aggregates features from windows of diverse shapes to calibrate and expand the attention area, involving more self-similar textures. Specifically, we propose a Distortion Modulated Rectangle-window Self-Attention (DMRSA) mechanism, which takes into account more self-similar regions of ERP images within one window. We integrate a deformable self-attention mechanism, which considers irregular neighborhoods and captures out-of-window similar patterns [9], with DMRSA. Self-similarity is essential because it significantly contributes to the reconstruction of HR images [10, 11, 12, 13, 14]. In DMRSA, Rectangle-window Self-Attention (Rwin-SA) is modulated by the distortion guidance that is generated through a newly devised Distortion Guidance Generator (DGG). DGG considers the variability of geometric distortion in distinct latitudes [1]. Furthermore, we dynamically aggregate the features from two self-attention modules by exploiting the information of windows with various shapes in Dynamic Feature Aggregation (DFA). To reveal the effects on the range of the involved area, we resort to Local Attribution Map (LAM) [5], which is an attribution method for analyzing and visualizing attribution in SR. As shown in Fig. 1, our proposed method can utilize a wider range of self-similar information and more pixels to recover the patch in the red box, achieving a higher Diffusion Index (DI) and better SR results.
@@
  • Based on the distortion property of ERP images, we propose a novel framework for omnidirectional image super-resolution, called GDGT-OSR. Our method perceives the distortion and adapts to the unevenly distorted content in ERP images, enabling it to associate with more self-similar textures in the reconstruction process.
  • We propose a Distortion Modulated Rectangle-window Self-Attention (DMRSA) mechanism, integrated with deformable self-attention. These two distinct self-attention mechanisms can involve more self-similar and related features from windows of various shapes, which are more appropriate for capturing similar patches in ERP images.
  • We propose a Distortion Guidance Generator (DGG) to transform the geometric distortion into distortion guidance, exploiting the variability of distortion across latitudes. To better adapt to the distorted content in ERP images, the distortion guidance is utilized to modulate the Rectangle-window Self-Attention (Rwin-SA) mechanism.
  • We propose a Dynamic Feature Aggregation (DFA) module to dynamically aggregate the features based on the differences between them. This can complement and thus better integrate the features from different self-attention mechanisms.
  • Our GDGT-OSR framework achieves superior performance compared with other state-of-the-art methods, both quantitatively and qualitatively.
The remaining parts of this paper are organized as follows. In Section II, we briefly review some related works. Section III firstly introduces preliminaries about the related knowledge, followed by a detailed introduction of the proposed GDGT-OSR framework. Section IV presents the experimental settings, experiment results, and analysis of ablation studies. Finally, Section V provides a brief conclusion of this paper.
@@ II-A Single Image Super-Resolution
Long before the emergence of deep learning, researchers developed learning-based image resolution enhancement methods using image pyramids [15] and look-up tables [16, 17, 18]. After deep learning models demonstrated powerful capabilities in many related applications, researchers began to apply deep learning models, such as convolutional neural networks (CNNs), to image SR [19]. As recent literature is dominated by deep-learning-based methods, we review three types of popular deep-learning-based SR methods, including GAN-based SR, Transformer-based SR, and Diffusion-based SR.
@@
Since 2017, Transformers [27] have been widely used in computer vision. Researchers have applied Transformers to SR and demonstrated their efficacy in low-level computer vision tasks [28, 7, 8]. IPT [28] introduces a backbone model based on the standard Transformer to address diverse restoration challenges. Based on Swin Transformer [29], SwinIR [7] is proposed to solve different image restoration problems and has become a strong and popular backbone architecture for image restoration. However, using local squared windows, SwinIR suffers from a lack of direct interaction among windows, restricting the potential of establishing long-range dependencies. To solve this issue, CAT [8] proposes the Rwin-SA mechanism to broaden the attention area and increase the interaction across various windows by utilizing horizontal and vertical rectangle window attention.
@@ II-B Omnidirectional Image Super-Resolution
Deng et al. [1] pointed out that the pixel density is non-uniform and varies across latitudes in ERP projected ODIs. Therefore, they propose a progressive pyramid network, namely LAU-Net, to super-resolve the pixels at different latitude bands hierarchically, rather than making the model adapt to ODIs. However, training multiple levels of networks for different latitude bands is computationally expensive, and it can also lead to inconsistencies between latitude bands. The work [44] directly concatenates a distortion map and an LR image as the input, which is fed into a network designed for 2D plain images. However, the information of the distortion map is not exploited sufficiently, and it is difficult for a 2D-image SR network to learn to restore ODIs without modification. SphereSR [45] generates a continuous spherical image representation and employs the local implicit image function (LIIF) [46] to predict the RGB values under various projection types continuously. Although SphereSR can resolve ODIs with arbitrary projection types flexibly, it requires training multiple network branches for different projection types. When generating training pairs, all of these ODISR methods apply uniform bicubic downsampling on the ERP images, neglecting the geometric properties of ERP in the degradation process. OSRT [47] utilizes Fisheye downsampling that applies uniform bicubic downsampling on the original ODIs. Even though OSRT makes use of the deformable Transformer to leverage the distortion information, the attention area is still small and inefficient for reconstructing ODIs and it is difficult to adaptively calibrate the features across latitudes. To solve this issue, we propose to enlarge the attention area by involving different kinds of self-attention with diverse window shapes. Furthermore, we introduce a distortion guidance generator to modulate the features based on the variability of distortion across latitudes.
@@ III-C Loss Function
where $I^{gt}$ and $I^{o}$ denote the ground-truth image and the output image, respectively. $D(h,w)$ represents the value of the distortion map at the coordinate $(h,w)$. From Eq. (6), we can know that the weights are larger at lower altitudes, while the weights are smaller at higher altitudes. This design is reasonable from the perspective of non-uniform pixel density: the pixel density is more compact at lower altitudes, which means more pixels are distributed around lower altitudes, so these pixels should contribute more to the loss function, and vice versa.
@@
TABLE II: Ablation study of DMRSA and DDSA. The scaling factor is 4.
@@
TABLE III: Ablation study of DGG and Diff. The scaling factors are 2 and 4.
@@
TABLE IV: Result comparison among variants of DGG. The scaling factor is 4.
@@
IV Experiments
IV-A Experimental Settings
IV-B Comparisons with State-of-the-Art Methods
Quantitative Results. Table I shows the quantitative comparisons among different methods under ×$\times 2$ and ×$\times 4$ scaling factors. OSRT is an ODI-SR method, and we report the results as in [47]. The remaining compared methods, e.g., RCAN, SRResNet, EDSR, and SwinIR, are SR methods for 2D plain images. Thus, we retrained them on the ODI-SR dataset plus the auxiliary DF2K-ERP dataset. RCAN, SRResNet, EDSR, and SwinIR have inferior performance to OSRT and our GDGT-OSR, which demonstrates that the key to ODISR is the adaptiveness of the distortion information. With the dual-attention aggregation and distortion-guided components, our GDGT-OSR has strong distortion-awareness and distortion transformation abilities and achieves state-of-the-art (SOTA) performance on the two public datasets for both ×$\times 2$ and ×$\times 4$ tasks. This illustrates that our proposed method can better exploit and transform the distortion map to enhance the SR performance than OSRT.
Qualitative Results.
Figure 9: Visual comparison among different methods on the ODI-SR dataset[1]. The scaling factor is 4.
@@ IV-B Comparisons with State-of-the-Art Methods
TABLE V: Result comparison among different ways of aggregation on SUN 360 PANORAMA [59]. The scaling factor is 4.
@@
Fig. 9 shows the visual comparison among different SR methods on the ODI-SR dataset, under a scaling factor of 4. The patches in red boxes are zoomed in for better visualization. Even though the SR methods for 2D plain images are retrained with the ODI datasets, their performances are still worse than those of OSRT and our GDGT-OSR. For example, as shown in the third sample in Fig. 9, the ODISR methods, i.e., OSRT and our GDGT-OSR, can restore more visually pleasing results with less blurriness and distortion of the chair than the preceding four methods. This demonstrates that the inclusion and exploration of the distortion map contribute to the superiority and effectiveness of the last two ODISR methods. Although OSRT represents SOTA performance in ODISR methods, our proposed method surpasses it by restoring more details and richer textures. For example, in the last sample in Fig. 9, our method can restore comparatively complete window frames, while some parts of the window frames are missing in the visual result of OSRT.
TABLE VI: Ablation study of the window size of DMRSA. The scaling factor is 4.
@@
TABLE VII: Result comparison between the utilization of $l1$ loss and WS-$l1$ loss as the training loss. The scaling factor is 4.
@@
IV-C Ablation Study
Impacts of Self-Attentions. We investigate the influence of DMRSA and DDSA. As shown in Table II, compared to the variant without both DMRSA and DDSA, the performance of the variant with DMRSA improves significantly. When both DMRSA and DDSA are incorporated into the network architecture, the performance is further improved, which demonstrates the effectiveness of the proposed GDGT-OSR. Moreover, we investigate the impacts of the offsets in DDSA. Fig. 10 shows the visual comparison of offset maps in DDSAs of both OSRT [47] and GDGT-OSR. We can see that the offsets from our GDGT-OSR are larger than those from OSRT around the polar area where the distortion is the most severe, while the offsets from both OSRT and GDGT-OSR around the equator area are small because of negligible distortion. This observation demonstrates that, when combined with DMRSA, DDSA can adapt better to the distortion and further broaden its attention areas to involve a larger range of pixels. The above results and analysis show the effectiveness of collaboration between DMRSA and DDSA.
Impacts of Distortion Guidance Generator. DGG is designed to thoroughly exploit the modulation information of the distortion map for Rwin-SA. As shown in Table III, we present the results of our model without and with DGG in the second and fourth rows under both ×$\times 2$ and ×$\times 4$ scaling factors. With DGG, the model performance surpasses its counterpart (without DGG) across all PSNR-related and SSIM-related metrics, demonstrating the effectiveness of DGG. Fig. 11 shows the qualitative comparison of our model without and with DGG. As shown in the red cropped patches, while the visual results of the model without DGG differ from the ground truth and appear distorted, the model with DGG can restore more visually pleasing results. It demonstrates that our model benefits from DGG in enhancing SR performances.
To further investigate the effects of DGG, we conducted some ablation experiments on the architecture of DGG, as shown in Table IV, i.e., different ways of exploiting the distortion map. ‘Direct’ represents that the distortion map is replicated along the channel dimension, and then directly utilized to modulate the feature maps. ‘One conv’ means that we use one convolutional layer to replace the proposed DGG. ‘w/o atten’ denotes the DGG without the attention branch. It can be seen that the performance deteriorates most in the ‘direct’ variant. Although we use one convolutional layer to encode the distortion map, the performance is almost the same as that of the ‘direct’ variant. Thus, we can conclude that exploiting the distortion map in simple ways brings little improvement. Moreover, the performance of the ‘w/o atten’ variant is inferior to that of GDGT-OSR, which involves a complete DGG. DGG not only encodes the distortion map into feature maps that match the dimension of the key/value features in Rwin-SA, but also effectively exploits the information of the distortion map.
TABLE VIII: Comparison of average DI values at four different pixels in the SUN 360 Panorama dataset[59].

### v2

@@ Abstract
As virtual and augmented reality applications gain popularity, omnidirectional image (ODI) super-resolution has become increasingly important. Unlike 2D plain images that are formed on a plane, ODIs are projected onto spherical surfaces. Applying established image super-resolution methods to ODIs, therefore, requires performing equirectangular projection (ERP) to map the ODIs onto a plane. ODI super-resolution needs to take into account geometric distortion resulting from ERP. However, without considering such geometric distortion of ERP images, previous methods only utilize a limited range of pixels and may easily miss self-similar textures for reconstruction. In this paper, we introduce a novel Geometric Distortion Guided Transformer for Omnidirectional image Super-Resolution (GDGT-OSR). Specifically, a distortion modulated rectangle-window self-attention mechanism, integrated with deformable self-attention, is proposed to better perceive the distortion and thus involve more self-similar textures. Distortion modulation is achieved through a newly devised distortion guidance generator that produces guidance for the rectangular windows by exploiting the variability of distortion across latitudes. Furthermore, we propose a dynamic feature aggregation scheme to adaptively fuse the features from different self-attention modules. We present extensive experimental results on public datasets and show that the new GDGT-OSR outperforms methods in existing literature.
@@
publicationid: pubid: Copyright © 2024 IEEE. Personal use of this material is permitted. However, permission to use this material for any other purposes must be obtained from the IEEE by sending an email to pubs-permissions@ieee.org.
@@ I Introduction
One way to tackle this issue is through image super-resolution (SR), which reconstructs a high-resolution (HR) image from a low-resolution (LR) input [2, 3, 4]. For convenience in storage and transmission, raw ODIs are generally projected into 2D planar representations. Equirectangular projection (ERP) is the most common method for representing ODIs [5].
@@
Transformer has emerged as a powerful and versatile computational paradigm [7]. Local square-window self-attention is proposed to reduce the computational complexity of the global self-attention mechanism in vision transformer, leading to a limitation in the receptive field [8, 9]. However, for omnidirectional image super-resolution (ODISR), the ERP expands the ODIs and introduces distortion into ERP images. For example, a circle in the ODI is distorted into an oval in the ERP image. Small squared windows struggle to capture the whole oval, while large squared windows may encompass irrelevant patterns. Therefore, square-window self-attention is a suboptimal option for reconstructing ERP images. Rectangular windows [9] can calibrate and expand the receptive field by involving more self-similar textures along the direction of stretching distortion, which is more appropriate for modeling features of ERP images than squared windows. However, existing ODISR methods [1, 10, 11] fail to consider extracting features of distorted ERP images from the perspective of the window’s shape, leading to the limitation of involved pixels and self-similar textures. To solve this issue, we propose a Geometric Distortion Guided Transformer for Omnidirectional image Super-Resolution (GDGT-OSR), which aggregates features from windows of diverse shapes to calibrate and expand the attention area, involving more self-similar textures. Specifically, we propose a Distortion Modulated Rectangle-window Self-Attention (DMRSA) mechanism, which takes into account more self-similar regions of ERP images within a rectangular window. We integrate a deformable self-attention mechanism, which considers irregular neighborhoods and captures out-of-window similar patterns [12], with DMRSA. Self-similarity is essential because it significantly contributes to the reconstruction of HR images [13, 14, 15, 16]. In DMRSA, Rectangle-window Self-Attention (Rwin-SA) is modulated by the distortion guidance that is generated through a newly devised Distortion Guidance Generator (DGG). DGG transforms the geometric distortion into distortion guidance, making the DMRSA adapt to the stretched ERP images. Furthermore, we dynamically aggregate the features from two self-attention modules by exploiting the information of windows with various shapes in Dynamic Feature Aggregation (DFA). To reveal the effects on the range of the involved area, we resort to Local Attribution Map (LAM) [6], which is an attribution method for analyzing and visualizing attribution in SR. As shown in Fig. 1, our proposed method can utilize a wider range of self-similar information and more pixels to recover the patch in the red box, achieving a higher Diffusion Index (DI) and better SR results.
@@
  • We propose a novel framework, GDGT-OSR, designed for omnidirectional image super-resolution. By leveraging the distortion as guidance, GDGT-OSR effectively captures self-similar textures with windows of different shapes, leading to enhanced SR performance.
  • To leverage the distortion information, we propose a Distortion Modulated Rectangle-window Self-Attention (DMRSA) mechanism paired with a Distortion Guidance Generator (DGG). The DGG transforms geometric distortion into distortion guidance, empowering DMRSA to adapt to the inherent distortion in ERP images through the newly designed distortion-guided rectangular windows.
  • We devise a Dynamic Feature Aggregation (DFA) module to aggregate and complement features from different self-attention mechanisms based on the differences between them.
  • Our GDGT-OSR framework achieves superior performance in ODISR, outperforming other state-of-the-art methods, including 2D plain SR and ODISR methods, in terms of quantitative and qualitative results.
The remaining parts of this paper are organized as follows. In Section II, we briefly review some related works. Section III introduces preliminaries about the related knowledge, followed by a detailed introduction of the proposed GDGT-OSR framework. Section IV presents the experimental settings, experiment results, and analysis of ablation studies. Finally, Section V-B provides a summary, including the limitations and a brief conclusion of this paper.
@@ II-A Single Image Super-Resolution
Long before the emergence of deep learning, researchers developed learning-based image resolution enhancement methods using image pyramids [17] and look-up tables [18, 19, 20]. After deep learning models demonstrated powerful capabilities in many related applications, researchers began to apply deep learning models, such as convolutional neural networks (CNNs), to image SR [21, 2, 3, 4]. As recent literature is dominated by deep-learning-based methods, we review three types of popular deep-learning-based SR methods, including GAN-based SR, Transformer-based SR, and Diffusion-based SR.
@@
Since 2017, Transformers [29] have been widely used in computer vision. Researchers have applied Transformers to SR and demonstrated their efficacy in low-level computer vision tasks [30, 8, 9, 31, 32]. IPT [30] introduces a backbone model based on the standard Transformer to address diverse restoration challenges. Based on Swin Transformer [33], SwinIR [8] is proposed to solve different image restoration problems and has become a strong and popular backbone architecture for image restoration. However, using local squared windows, SwinIR suffers from a lack of direct interaction among windows, restricting the potential of establishing long-range dependencies. To solve this issue, CAT [9] proposes the Rwin-SA mechanism to broaden the attention area and increase the interaction across various windows by utilizing horizontal and vertical rectangle window attention.
@@ II-B Omnidirectional Image Super-Resolution
Deng et al. [1] pointed out that the pixel density is non-uniform and varies across latitudes in ERP projected ODIs. Therefore, they propose a progressive pyramid network, namely LAU-Net, to super-resolve the pixels at different latitude bands hierarchically, rather than making the model adapt to ODIs. However, training multiple levels of networks for different latitude bands is computationally expensive, and it can also lead to inconsistencies between latitude bands. The work [48] directly concatenates a distortion map and an LR image as the input, which is fed into a network designed for 2D plain images. However, the information of the distortion map is not exploited sufficiently, and it is difficult for a 2D-image SR network to learn to restore ODIs without modification. SphereSR [10] generates a continuous spherical image representation and employs the local implicit image function (LIIF) [49] to predict the RGB values under various projection types continuously. Although SphereSR can resolve ODIs with arbitrary projection types flexibly, it requires training multiple network branches for different projection types. TCCL-Net[50] proposes a Transformer and convolution collaborative learning network for ODISR, which extracts both long-range and short-range dependencies in an end-to-end manner. Furthermore, in the work [51], the same authors further explore the SR of panoramic images from the left-right view by combining binocular information and panoramic characteristics. When generating training pairs, all of these ODISR methods apply uniform bicubic downsampling on the ERP images, neglecting the geometric properties of ERP in the degradation process. OSRT [11] utilizes Fisheye downsampling that applies uniform bicubic downsampling on the original ODIs. Even though OSRT makes use of the deformable Transformer to leverage the distortion information, the attention area is still small and inefficient for reconstructing ODIs and it is difficult to adaptively calibrate the features across latitudes. To solve this issue, we propose to enlarge the attention area by involving different kinds of self-attention with diverse window shapes. Furthermore, we introduce a distortion guidance generator to modulate the features based on the variability of distortion across latitudes.
@@ III-C Loss Function
where $I^{gt}$ and $I^{o}$ denote the ground-truth image and the output image, respectively. $D(h,w)$ represents the value of the distortion map at the coordinate $(h,w)$. From Eq. (6), we can know that the weights are larger at lower latitudes, while the weights are smaller at higher latitudes. From the perspective of non-uniform pixel density, this design is reasonable: the pixel density is more compact at lower latitudes while it is sparser at higher latitudes, which means more pixels are distributed around lower latitudes, so these pixels should contribute more to the loss function, and vice versa.
@@
TABLE II: Quantitative comparison of different methods under large scaling factors, i.e., specifically ×$\times 8$ and ×$\times 16$. The best results are highlighted in bold.
Method Scale ODI-SR[1] SUN 360 Panorama[63]
PSNR SSIM WS-PSNR WS-SSIM PSNR SSIM WS-PSNR WS-SSIM
Bicubic ×$\times{8}$ 24.24 0.6525 23.51 0.6204 23.73 0.6369 23.22 0.6324
SwinIR[8] 25.09 0.6873 24.38 0.6619 24.97 0.6881 24.57 0.6953
HAT[66, 31] 25.15 0.6909 24.42 0.6653 25.12 0.6936 24.69 0.7008
RGT[32] 24.86 0.6889 24.08 0.6632 24.68 0.6893 24.23 0.6963
OSRT[11] 25.28 0.6929 24.53 0.6664 25.32 0.6964 24.86 0.7027
GDGT-OSR (ours) 25.32 0.6949 24.60 0.6687 25.42 0.6994 25.00 0.7068
Bicubic ×$\times{16}$ 22.66 0.6171 21.90 0.5785 22.04 0.6025 21.42 0.5899
SwinIR[8] 23.33 0.6388 22.58 0.6046 22.86 0.6292 22.28 0.6230
HAT[66, 31] 23.32 0.6401 22.55 0.6057 22.87 0.6306 22.27 0.6246
RGT[32] 23.17 0.6391 22.40 0.6049 22.70 0.6297 22.10 0.6236
OSRT[11] 23.49 0.6421 22.71 0.6074 23.11 0.6339 22.48 0.6278
GDGT-OSR (ours) 23.54 0.6425 22.78 0.6087 23.21 0.6356 22.60 0.6303
IV Experiments
IV-A Experimental Settings
IV-B Comparisons with State-of-the-Art Methods
Quantitative Results. Table I shows the quantitative comparisons among different methods under ×$\times 2$ and ×$\times 4$ scaling factors. OSRT is an ODI-SR method, and we report the results as in [11]. The remaining compared methods, e.g., RCAN [64], SRResNet [24], EDSR [65], SwinIR [8], HAT [31] and RGT [32] are SR methods for 2D plain images. Thus, we retrained them on the ODI-SR dataset plus the auxiliary DF2K-ERP dataset. RCAN, SRResNet, EDSR, and SwinIR have inferior performance to OSRT and our GDGT-OSR, which demonstrates that the key to ODISR is the adaptiveness of the distortion information. With the dual-attention aggregation and distortion-guided components, our GDGT-OSR has strong distortion-awareness and distortion transformation abilities and achieves state-of-the-art (SOTA) performance on the two public datasets for both ×$\times 2$ and ×$\times 4$ tasks. This illustrates that our proposed method can better exploit and transform the distortion map to enhance the SR performance than OSRT.
To further evaluate the effectiveness of GDGT-OSR on larger scaling factors, we conducted experiments with scaling factors of ×$\times 8$ and ×$\times 16$. The results, presented in Table II, compare GDGT-OSR with other SOTA SR methods, including SwinIR [8], HAT [31], RGT [32], and OSRT [11]. As shown in Table II, despite the increased difficulty associated with larger scaling factors, our proposed GDGT-OSR consistently outperforms other SOTA 2D plain image SR and ODISR methods. This demonstrates the robustness and generalization capability of GDGT-OSR across different scaling factors, highlighting its effectiveness.
Qualitative Results.
Figure 9: Visual comparison of different methods on the ODI-SR dataset[1] with a scaling factor of 4. The red boxes highlight zoomed-in patches for better visualization.
Figure 10: Visual comparison of different methods on the ODI-SR dataset[1] with a scaling factor of 8. The red boxes highlight zoomed-in patches for better visualization.
Fig. 9 presents a visual comparison of different SR methods on the ODI-SR dataset with a scaling factor of 4. The red-boxed patches are zoomed in for better visualization. Although the 2D plain image SR methods are retrained with the ODI datasets, their performances are still worse than those of OSRT and our GDGT-OSR. For example, as shown in the third sample in Fig. 9, the ODISR methods, i.e., OSRT and our GDGT-OSR, can restore more visually pleasing results with less blurriness and distortion of the chair compared to the preceding four methods. This demonstrates that the inclusion and exploration of the distortion map contribute to the superiority and effectiveness of the last two ODISR methods. Although OSRT represents SOTA performance in ODISR methods, our proposed method surpasses it by restoring more details and richer textures. For example, in the last sample in Fig. 9, our method can restore comparatively complete window frames, while some parts of the window frames are missing in the visual result of OSRT.
Fig. 10 shows the visual comparison of different methods under the ×$\times 8$ scaling factor. It is more difficult to super-resolve the images under a larger scaling factor because more details are lost in the LR images. As shown in Fig. 10, under the ×$\times 8$ scaling factor, most of the SOTA SR methods, including SwinIR [8], HAT [31], RGT [32] and OSRT [11], struggle to recover the details and textures, leading to unsatisfactory visual results. With the proposed dual-attention and distortion-guided mechanisms, GDGT-OSR effectively manages distortion transformation and can recover more details for the ERP images even under the challenging large scaling factor. To summarize, the above visual results demonstrate the superiority and effectiveness of the proposed GDGT-OSR under various scaling factors.
TABLE III: Ablation study of DMRSA and DDSA. The scaling factor is 4.
@@ IV-B Comparisons with State-of-the-Art Methods
TABLE IV: Ablation study of DGG and Diff. The scaling factors are 2 and 4.
@@
TABLE V: Result comparison among variants of DGG. The scaling factor is 4.
@@
IV-C Ablation Study
Impacts of Self-Attentions. We investigate the influence of DMRSA and DDSA. As shown in Table III, compared to the variant without both DMRSA and DDSA, the performance of the variant with DMRSA improves significantly. The variant with only DMRSA performs better than that with only DDSA. This illustrates that the proposed DMRSA mechanism dominantly contributes to the ODISR. When both DMRSA and DDSA are incorporated into the network architecture, the performance is further improved, which demonstrates the effectiveness of the proposed GDGT-OSR framework. Moreover, we investigate the impacts of the offsets in DDSA. Fig. 11 shows the visual comparison of offset maps in DDSAs of both OSRT [11] and GDGT-OSR. We can see that the offsets from our GDGT-OSR are larger than those from OSRT around the polar area where the distortion is the most severe, while the offsets from both OSRT and GDGT-OSR around the equator area are small due to negligible distortion. This observation demonstrates that, when combined with DMRSA, DDSA can adapt better to the distortion and further broaden its attention areas to involve a larger range of pixels. The above results and analysis show the effectiveness of collaboration between DMRSA and DDSA.
@@ IV-C Ablation Study
Impacts of Distortion Guidance Generator. DGG is designed to thoroughly exploit the modulation information of the distortion map for Rwin-SA. As shown in Table IV, we present the results of our model without and with DGG in the second and fourth rows under both ×$\times 2$ and ×$\times 4$ scaling factors. With DGG, the model performance surpasses its counterpart (without DGG) across all PSNR-related and SSIM-related metrics, demonstrating the effectiveness of DGG. Fig. 12 shows the qualitative comparison of our model without and with DGG. As shown in the red cropped patches, while the visual results of the model without DGG differ from the ground truth and appear distorted, the model with DGG can restore more visually pleasing results. It demonstrates that our model benefits from DGG in enhancing SR performances.
To further investigate the effects of DGG, we conducted some ablation experiments on the architecture of DGG, as shown in Table V, i.e., different ways of exploiting the distortion map. ‘Direct’ represents that the distortion map is replicated along the channel dimension, and then directly utilized to modulate the feature maps. ‘One conv’ means that we use one convolutional layer to replace the proposed DGG. ‘w/o atten’ denotes the DGG without the attention branch. It can be seen that the performance deteriorates most in the ‘direct’ variant. Although we use one convolutional layer to encode the distortion map, the performance is almost the same as that of the ‘direct’ variant. Thus, we can conclude that exploiting the distortion map in simple ways brings little improvement. Moreover, the performance of the ‘w/o atten’ variant is inferior to that of GDGT-OSR, which involves a complete DGG. DGG not only encodes the distortion map into feature maps that match the dimension of the key/value features in Rwin-SA, but also effectively exploits the information of the distortion map.
Impacts of Diff. In DFA, we calculate the difference between $F^{DMRSA}$ and $F^{DDSA}$, which is used to adaptively weigh the combined features. As shown in Table IV, the performance slightly degrades without Diff. Fig. 13 shows the visual comparison between our model without and with the Diff design in DFA. We can see that some details are missing without Diff, while more details are restored with Diff. It is worth noting that computing the Diff does not require additional parameters, but it can bring improvements in restoring details. In Table IV, we can observe that the SR performance is moderately impacted when both DGG and Diff are absent simultaneously, especially under the ×$\times 2$ scaling factor. This demonstrates the necessity and importance of both utilizing the distortion map and leveraging the distinction between features from DMRSA and DDSA.
@@
Impacts of Dynamic Feature Aggregation. We propose to aggregate features from DMRSA and DDSA in a dynamic way, based on the learned weights. To verify the efficacy of DFA, we compare the SR performance of different aggregation methods on the SUN 360 panorama dataset, as illustrated in Table VI. ‘Addition’ denotes that the two features are summed directly to obtain the final feature. ‘Concatenation’ denotes that the two features are concatenated along the channel dimension, and then convolved by a convolutional layer to reduce the number of channels to match that of the input features. From Table VI, we can see that DFA achieves the best or comparative performance across all evaluation metrics compared with other aggregation methods.
TABLE VI: Result comparison among different ways of aggregation on SUN 360 PANORAMA [63]. The scaling factor is 4.
@@
Impacts of Window Size. We investigate the effects of different sizes of rectangle windows, as shown in Table VII. A window size of (32/8) means that the height and width are 32 and 8 for vertical windows, respectively, and 8 and 32 for horizontal windows. As demonstrated in Table VII, the network with a window size of (64/8) outperforms those with window sizes of (32/8) and (64/4). This improvement may be because the rectangle window with size (64/8) can effectively capture features of distorted content in a more appropriate range of area.
Impacts of Window Shape. Besides the window size, the effects of the window shapes of the rectangular windows in DMRSA are also worth exploring. To this end, we conducted experiments using various configurations of rectangular window orientations, i.e., vertical only, horizontal only, and a combination of both vertical and horizontal windows with 50% each. The results, as presented in Table VIII, indicate that the configuration with only horizontal rectangular windows outperforms the one with only vertical rectangular windows across both testing datasets. This advantage is likely attributable to the substantial horizontal stretch of projected ERP images, enabling horizontal rectangular windows to capture more self-similar features for enhanced reconstruction. Nonetheless, as demonstrated in Table VIII, the variant incorporating both vertical and horizontal rectangular windows achieves superior performance. Despite the primary horizontal distortion of ERP images, certain self-similar textures and patterns may still exist vertically, which can be effectively captured using vertical windows. Consequently, our DMRSA mechanism employs a combination of both vertical and horizontal rectangular windows.
TABLE VII: Ablation study of the window size of DMRSA. The scaling factor is 4.
@@
TABLE VIII: Ablation study of the window shape of DMRSA. The scaling factor is 4.
Window shape ODI-SR[1] SUN 360 Panorama[63]
PSNR SSIM WS-PSNR WS-SSIM PSNR SSIM WS-PSNR WS-SSIM
Vertical 27.40 0.7773 26.68 0.7619 27.83 0.7842 27.74 0.8027
Horizontal 27.42 0.7771 26.70 0.7616 27.86 0.7842 27.76 0.8026
Vertical & Horizontal 27.44 0.7776 26.72 0.7623 27.92 0.7850 27.81 0.8036
TABLE IX: Result comparison among different training losses. The scaling factor is 4.
@@
Figure 14: Visual comparison of local attribution maps and SR results among different methods. The Diffusion Index (DI) is given below the local attribution maps. A higher DI indicates a wider range of the involved pixels, and vice versa.
TABLE X: Model complexity comparison. The scaling factor is 8.
Model #Params(M) #Multi-Adds(G) PSNR WS-PSNR
SwinIR 12.05 60.59 25.09 24.38
HAT 20.92 96.05 25.15 24.42
RGT 13.51 62.17 24.86 24.08
OSRT 12.08 54.50 25.28 24.53
GDGT-OSR 15.99 69.74 25.32 24.60
TABLE XI: Comparison of average DI values at four different pixels in the SUN 360 Panorama dataset[63].